# The most symmetric smooth cubic surface over a finite field of characteristic $2$

### v1

@@ 1. Introduction
was interesting to many mathematicians since ancient times, at least since Porisms of Diophantus, where he mentioned that the difference of two cubes can be represented as a sum of two cubes (see [4, Chapter \@slowromancapxxi@] for more historical details). It was studied a lot in number theory and algebraic geometry. The Fermat cubic surface has the largest automorphism group among smooth cubic surfaces over an algebraically closed field of characteristic not equal to $3$ This was proved by T. Hosoh in [8, Theorem 5.3] for fields of characteristic zero and, by I. Dolgachev and A. Duncan in [6, Theorem 1.1] for fields of positive characteristics. In particular, we have the following theorem:
Theorem 1.1 ([6, Theorem 1.1]).
@@
Note that from the proof of Theorem 1.1 (see [6, §5.1] for more details) we get the following theorem about automorphism group of the Fermat cubic surface over $\mathbb{F}_{4}.$
Theorem 1.2 ([6, §5.1]).
@@
Theorem 1.3 ([13, Theorem 3 and Example 9]).
@@
The following proposition shows that over fields $\mathbb{F}_{2^{2k+1}}$ the Fermat cubic surface and the cubic (1.1) are not isomorphic whereas over $\overline{\mathbb{F}}_{2}$ they are isomorphic.
@@
Let $\mathcal{F}$ be the Fermat cubic surface over a field $\mathbf{F}$ of characteristic $2$ Then if $\mathbf{F}=\mathbb{F}_{2^{2k+1}}$ we get $\mathrm{Aut}(\mathcal{F})\simeq\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}.$
@@
Now let us describe the structure of the paper. In Section 2 we study subgroups of order $5$ in the automorphism group of $\mathbb{P}^{1}$ over finite fields of characteristic $2$ In Section 3 we study quadric surfaces and orbits of points on quadric surfaces under the action of the group of order $5$ In Section 4 we study elements and subgroups of Weyl group $W(\mathrm{E}_{6}).$ In Section 5 we recall some basic facts about smooth cubic surfaces and discuss their automorphisms. In Section 6 we prove Theorems 1.4 and 1.5. In Section 7 we discuss the Fermat cubic surface over finite fields of characteristic $2$ and prove Proposition 1.6 and Corollary 1.7.
@@
Acknowledgment. It is a pleasure for the author to express gratitude to C. A. Shramov for his care, for proposing this problem, a lot of discussion and constant attention. Also the author would like to thank A. S. Trepalin who taught the author his vision of cubic surfaces and their symmetries.
This work was performed at the Steklov International Mathematical Center and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2022-265). The author is a winner of the all-Russia mathematical August Moebius contest of graduate and undergraduate student papers and thanks the jury and the board for the high praise of his work. The author was partially supported by Theoretical Physics and Mathematics Advancement Foundation “BASIS” and by the HSE University Basic Research Program.
@@ 2. Projective line
Similarly, assertion (ii) follows from the equations
$(\xi+\xi^{4})^{2^{2k+1}}=\xi^{2\cdot 4^{k}}+\xi^{8\cdot 4^{k}}=\xi^{2\cdot(-1)^{k}}+\xi^{8\cdot(-1)^{k+1}}=\xi^{2}+\xi^{3}$
$(\xi^{2}+\xi^{3})^{2^{2k+1}}=\xi^{4\cdot 4^{k}}+\xi^{6\cdot 4^{k}}=\xi^{(-1)^{k+1}}+\xi^{(-1)^{k}}=\xi+\xi^{4}.$
By definition, $\psi(x)=x^{2^{2k+1}}$ for $x\in\mathbb{F}_{4^{2k+1}}.$ Therefore, assertion (iii) immediately follows from assertions (i) and (ii).
@@
The order of this group $\mathrm{PGL}_{2}(\mathbb{F}_{2^{2k+1}})$ is equal to
@@
which is isomorphic to $\mathbb{Z}/5\mathbb{Z},$ there is a matrix $M\in\mathrm{GL}_{2}(\mathbf{F})$ of order $5$ such that $\rho(M)$ is a generator of $G.$ Since $\xi$ lies in $\mathbf{F}$ the matrix $M$ is diagonalizable. So up to conjugation and multiplication by $\xi^{r}$ for an integer $r$ we have $M=\bigl{(}\begin{smallmatrix}1&0\\ 0&\xi\end{smallmatrix}\bigr{)}.$ This means that the group of order $5$ in $\mathrm{PGL}_{2}(\mathbf{F})$ is unique up to conjugation.
@@
where the minimal polynomial of the first matrix is $x^{2}+(\xi+\xi^{4})x+1$ and the minimal polynomial of the second matrix is $x^{2}+(\xi^{2}+\xi^{3})x+1.$ Note that if $M$ is conjugate to $\bigl{(}\begin{smallmatrix}\xi+\xi^{4}&1\\ 1&0\end{smallmatrix}\bigl{)},$ then $M^{2}$ is conjugate to $\bigl{(}\begin{smallmatrix}\xi^{2}+\xi^{3}&1\\ 1&0\end{smallmatrix}\bigl{)}$ and vice versa. Thus, up to conjugation the group of order $5$ in $\mathrm{PGL}_{2}(\mathbb{F}_{4^{k}})$ can be generated by the matrix (2.2), and so up to conjugation such group is unique.
@@ 3. Quadric surfaces
In this section we study smooth quadric surfaces in $\mathbb{P}^{3}$ over finite fields of characteristic $2$ First of all, we refer the reader to [7, §8.1] or [14, Section 1.3] for the definition of Weil restriction of scalars. We recall the result from the paper [11] about the classification of quadric surfaces in $\mathbb{P}^{3}.$
Lemma 3.1 (cf. [11, Lemma 7.3]).
@@
[…]
@@
where $a$ and $b$ are non-zero and $\psi\in\mathrm{Gal}(\mathbb{F}_{4^{2k+1}}/\mathbb{F}_{2^{2k+1}})$, can be obtained from the point ([1:1],[1:1])([1:1],[1:1])( [ 1 : 1 ] , [ 1 : 1 ] ) by the action of the automorphism
@@ 4. The Weyl group of type $\mathrm{E}_{6}$
Lemma 4.2.
The subgroup $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ in $W(\mathrm{E}_{6})$ coincides with the subgroup of elements with determinant equal to $1$ considered as the action on the root system $\mathrm{E}_{6}.$
Proof.
As all elements in $W(\mathrm{E}_{6})$ are isometries their determinant is ±$\pm 1.$ So the subgroup of elements with determinant equal to $1$ is a normal subgroup of index $2$ since it is a kernel of the surjection $W(\mathrm{E}_{6})\twoheadrightarrow\mathbb{Z}/2\mathbb{Z}.$ Let $K$ be the kernel of this homomorphism. Then $H=\mathrm{PSU}_{4}(\mathbb{F}_{2})\cap K$ is a normal subgroup in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ By [15, §3.6.1] the group $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ is simple, thus $H$ is either trivial, or coincides with $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ However, $H$ is not trivial since $g^{2}\in\mathrm{PSU}_{4}(\mathbb{F}_{2})$ for any element $g\in K.$ Therefore, $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ coincides with $K.$
Now we recall the results from [1, Table 9] and [12, Theorem 1] about the conjugacy classes of the elements of order $2$, $5$ and $10$ in the Weyl group $W(\mathrm{E}_{6}).$
@@
As by Lemma 4.3 the order of the centralizer of an element $g$ of order $5$ in $W(\mathrm{E}_{6})$ is equal to $10$ This means that the centralizer of $g$ in the Weyl group is isomorphic either to $\mathbb{Z}/10\mathbb{Z},$ or to a dihedral group of order $10,$ since these two groups are the only groups of order $10$ up to isomorphism. However, the second case is impossible because $g$ lies in its centralizer.
@@ 5. Cubic surfaces
In this section we introduce the notation related to smooth cubic surfaces and recall the description of their geometry. More details on smooth cubic surfaces can be found in [5, Chapter 9] or [10, Chapter \@slowromancapiv@].
@@
$\Delta=\{\alpha\in\mid K_{S_{\overline{\mathbf{F}}}}\cdot\alpha=0\;\text{and}\;\alpha^{2}=-2\}.$
Then this set is the root system $\mathrm{E}_{6}.$ The Weyl group $W(\mathrm{E}_{6})$ acts on the Picard group $\mathrm{Pic}(S_{\overline{\mathbf{F}}})$ by orthogonal transformations with respect to the intersection form. It fixes $K_{S_{\overline{\mathbf{F}}}}$ and preserves the cone of effective divisors.
@@
Theorem 5.1 (see, for instance, [5, Corollary 8.2.40]).
@@
Using the classification of automorphism groups of smooth cubic surfaces we can say more about the group, where automorphism groups lie.
@@
Let $S$ be a smooth cubic surface. Then $\mathrm{Aut}(S)\subseteq\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
@@
This directly follows from the classification of the automorphism groups which can be found in [8, Theorem 5.3] in zero characteristic and in [6, Theorem 1.1] for positive characteristic.
@@
Lemma 5.3 ([3, Table 7.1, column 8]).
Let $S$ be a smooth cubic surface over an algebraically closed field. Consider an element $g\in W(\mathrm{E}_{6})$ of order $2$ acting on $S.$ The following assertions hold.
@@
    Assume that $g$ lies in the conjugacy class $A_{1}.$ Then there are $15$ lines on $S$ which are invariant under the action of $g.$ Moreover, there are $6$ pairs of skew lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$
@@
    Assume that $g$ lies in the conjugacy class $A_{1}^{2}.$ Then there are $7$ lines on $S$ which are invariant under the action of $g.$ There are $8$ pairs of skew lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$ Moreover, there are $2$ pairs of intersecting lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$
@@
    Assume that $g$ lies in the conjugacy class $A_{1}^{3}.$ Then there are $3$ lines which are invariant under the action of $g.$ There are $6$ pairs of skew lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$ Moreover, there are $6$ pairs of intersecting lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$
@@
    Assume that $g$ lies in the conjugacy class $A_{1}^{4}.$ Then there are $3$ lines which are invariant under the action of $g.$ Moreover, there are $12$ pairs of intersecting lines which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$
Remark 5.4.
Let $S$ be a smooth cubic surface over a field $\mathbf{F}.$ Note that the natural action of the group $G\simeq\mathfrak{S}_{6}$ on the Picard group of $S_{\overline{\mathbf{F}}}$ which permutes the indices of exceptional curves $E_{1},\ldots,E_{6}$ in not geometric, i.e. $G$ does not lie in the automorphism group of $S_{\overline{\mathbf{F}}}.$ Suppose the contrary. Let us consider the transposition $(12).$ This element fixes the following lines
(5.1) $L_{12},\,L_{34},\,L_{35},\,L_{36},\,L_{45},\,L_{46},\,L_{56},\,E_{3},\,E_{4},\,E_{5},\,E_{6},\,Q_{3},\,Q_{4},\,Q_{5},\,Q_{6}.$
therefore, by Lemma 5.3 the transposition lies in conjugacy class $A_{1}.$ By Lemma 4.3 its determinant is equal to $-1.$ So by Lemma 4.2 the transposition does not lie in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Remark 5.5.
Let us consider an element $\sigma$ of order $2$ in $W(\mathrm{E}_{6})$ which interchanges $E_{i}$ and $Q_{i}$ for all $i\in\{1,\ldots,6\}$ and fixes $L_{ij}$ for all $i,j\in\{1,\ldots,6\}$ and $i\neq j.$ Then $\sigma$ lies in the conjugacy class of type $A_{1}$ by Lemma 5.3 since $\sigma$ fixes $15$ lines.
@@
By [9, Chapter 4, Proposition 2.7 and Exercise 2.9] the blowup of points in general position on a quadric surface $\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}$ gives a smooth cubic surface. In other words, the preimages of points in general position on a quadric surface are the lines in quadratic position. Conversely, the image of the blowdown of lines in quadratic position on a smooth cubic surface is a set of points in general position on $\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}.$
@@
Let $S$ be a smooth cubic surface over an algebraically closed field $\mathbf{F}.$ Then the lines
(5.2) $L_{12},\;L_{13},\;L_{14},\;L_{15},\;L_{16}.$
are in quadratic position.
@@
By definition, the lines (5.2) are pairwise skew. Let us blow down the first four lines in (5.2). We obtain a morphism $f\colon S\to\widehat{S},$ where $\widehat{S}$ is a del Pezzo surface of degree $7$ and $f$ is the blowdown of $L_{12}$, $L_{13}$, $L_{14}$ and $L_{15}$. There are three $(-1)$-curves on $\widehat{S},$ namely, $f(L_{16}),$ $f(E_{6})$ and $f(Q_{6}).$ Moreover, we have the following intersection numbers:
$f(E_{6})\cdot f(Q_{6})=0\quad\text{and}\quad f(L_{16})\cdot f(E_{6})=f(L_{16})\cdot f(Q_{6})=1.$
So if we blow down $f(L_{16}),$ we get a del Pezzo surface of degree $8$ without $(-1)$-curves. This means that we get $\mathbb{P}^{1}\times\mathbb{P}^{1}.$
Let $S$ be a smooth cubic surface over a field $\mathbf{F}.$ Let us consider a Galois extension $\mathbf{L}$ of the field $\mathbf{F}.$ By [10, Theorem 23.8(ii)] the Galois group $\mathrm{Gal}(\mathbf{L}/\mathbf{F})$ acts on $\mathrm{Pic}(S_{\mathbf{L}})$ and preserves the canonical class of $S$ and the intersection form. This means that there is a homomorphism
$\mathrm{Gal}(\mathbf{L}/\mathbf{F})\to\mathrm{Aut}(\mathrm{Pic}(S_{\mathbf{L}}))\subset W(\mathrm{E}_{6}).$
Let $\Gamma$ be the image of the Galois group $\mathrm{Gal}(\mathbf{L}/\mathbf{F})$ in $W(\mathrm{E}_{6}).$ We remind the reader that $\Gamma$ lies in the centralizer of $\mathrm{Aut}(S),$ since any $\phi\in\mathrm{Aut}(S)$ is defined over $\mathbf{F}.$
@@
By Lemma 4.3 there is only one conjugacy class of order $5$ in $W(\mathrm{E}_{6}).$ Let us consider the element $g\in W(\mathrm{E}_{6})$ of order $5$ such that it acts on the indices of exceptional lines as the permutation $(23456).$ In particular, it permutes the lines
$L_{12},\;L_{13},\;L_{14},\;L_{15},\;L_{16}.$
By Lemma 5.8 these skew lines are in quadratic position. As up to conjugation the group $G$ is generated by $g,$ we get the desired.

### v2

@@ Abstract.
Keywords. Cubic surfaces, finite fields, automorphism groups.
Mathematics Subject Classification. 14J50
@@ 1. Introduction
was interesting to many mathematicians since ancient times, at least since Porisms of Diophantus, where he mentioned that the difference of two cubes can be represented as a sum of two cubes (see [5, Chapter \@slowromancapxxi@] for more historical details). It was studied a lot in number theory and algebraic geometry. The Fermat cubic surface has the largest automorphism group among smooth cubic surfaces over an algebraically closed field of characteristic not equal to $3$ This was proved by T. Hosoh in [9, Theorem 5.3] for fields of characteristic zero, and by I. Dolgachev and A. Duncan in [7, Theorem 1.1] for fields of positive characteristics. In particular, we have the following theorem:
Theorem 1.1 ([7, Theorem 1.1]).
@@
Note that from the proof of Theorem 1.1 (see [7, §5.1] for more details) we get the following theorem about automorphism group of the Fermat cubic surface over $\mathbb{F}_{4}.$
Theorem 1.2 ([7, §5.1]).
@@
Theorem 1.3 ([15, Theorem 3 and Example 9]).
@@
The following proposition shows that over fields $\mathbb{F}_{2^{2k+1}}$ the Fermat cubic surface and cubic (1.1) are not isomorphic whereas over $\overline{\mathbb{F}}_{2}$ they are isomorphic.
@@
Let $\mathcal{F}$ be the Fermat cubic surface over a field $\mathbf{F}$ of characteristic $2$ Then if $\mathbf{F}=\mathbb{F}_{2^{2k+1}},$ we get $\mathrm{Aut}(\mathcal{F})\simeq\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}.$
@@
Now let us describe the structure of the paper. In Section 2 we study subgroups of order $5$ in the automorphism group of $\mathbb{P}^{1}$ over finite fields of characteristic $2$ In Section 3 we study quadric surfaces and orbits of points on quadric surfaces under the action of the group of order $5$ In Section 4 we study elements and subgroups of the Weyl group $W(\mathrm{E}_{6}).$ In Section 5 we recall some basic facts about smooth cubic surfaces and discuss their automorphisms. In Section 6 we prove Theorems 1.4 and 1.5. In Section 7 we discuss the Fermat cubic surface over finite fields of characteristic $2$ and prove Proposition 1.6 and Corollary 1.7.
@@
Acknowledgment. It is a pleasure for the author to express gratitude to C. A. Shramov for his care, for proposing this problem, a lot of discussion and constant attention. Also the author would like to thank A. S. Trepalin who taught the author his vision of cubic surfaces and their symmetries. Also the author thanks the referee for important remarks about the geometric explanation of some group-theoretical lemmas.
This work was performed at the Steklov International Mathematical Center and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2022-265). The author is a winner of the all-Russia mathematical August Moebius contest of graduate and undergraduate student papers and thanks the jury and the board for the high praise of his work. The author was partially supported by Theoretical Physics and Mathematics Advancement Foundation “BASIS”.
@@ 2. Projective line
Assertions (ii) and (ii) follow from the fact that the Galois group of a finite extension of a finite field is generated by the Frobenius automorphism (see, for instance, [11, Chapter \@slowromancapv@, Theorem 5.4]).
@@
The order of the group $\mathrm{PGL}_{2}(\mathbb{F}_{2^{2k+1}})$ is equal to
@@
which is isomorphic to $\mathbb{Z}/5\mathbb{Z},$ there is a matrix $M\in\mathrm{GL}_{2}(\mathbf{F})$ of order $5$ such that $\rho(M)$ is a generator of $G.$ Since $\xi$ lies in $\mathbf{F}$ the matrix $M$ is diagonalizable. So up to conjugation and multiplication by $\xi^{r}$ for an integer $r$ we have $M^{s}=\bigl{(}\begin{smallmatrix}1&0\\ 0&\xi\end{smallmatrix}\bigr{)}$ for some $s=1,2,3,4.$ This means that the group of order $5$ in $\mathrm{PGL}_{2}(\mathbf{F})$ is unique up to conjugation.
@@
where the minimal polynomial of the first matrix is $x^{2}+(\xi+\xi^{4})x+1$ and the minimal polynomial of the second matrix is $x^{2}+(\xi^{2}+\xi^{3})x+1.$ Note that if $M$ is conjugate to $\bigl{(}\begin{smallmatrix}\xi+\xi^{4}&1\\ 1&0\end{smallmatrix}\bigl{)},$ then $M^{2}$ is conjugate to $\bigl{(}\begin{smallmatrix}\xi^{2}+\xi^{3}&1\\ 1&0\end{smallmatrix}\bigl{)},$ and vice versa. Thus, up to conjugation the group of order $5$ in $\mathrm{PGL}_{2}(\mathbb{F}_{4^{k}})$ can be generated by the matrix (2.2), and so up to conjugation such group is unique.
@@ 3. Quadric surfaces
In this section we study smooth quadric surfaces in $\mathbb{P}^{3}$ over finite fields of characteristic $2$ First of all, we refer the reader to [8, §8.1] or [16, Section 1.3] for the definition of the Weil restriction of scalars. We recall the result from the paper [13] about the classification of quadric surfaces in $\mathbb{P}^{3}.$
Lemma 3.1 (cf. [13, Lemma 7.3]).
@@
where $\mathrm{Aut}^{\circ}(X)\subset\mathrm{Aut}(X)$ is a subgroup that consists of all automorphism of $X$ which act trivially on $\mathrm{Pic}(X_{\mathbf{F}^{\text{sep}}}).$ More precisely, we have
@@
where $a$ and $b$ are non-zero and $\psi\in\mathrm{Gal}(\mathbb{F}_{4^{2k+1}}/\mathbb{F}_{2^{2k+1}})$ is a non-trivial element, can be obtained from the point ([1:1],[1:1])([1:1],[1:1])( [ 1 : 1 ] , [ 1 : 1 ] ) by the action of the automorphism
@@ 4. The Weyl group of type $\mathrm{E}_{6}$
Now we recall the results from [1, Table 9] and [14, Theorem 1] about the conjugacy classes of the elements of order $2$, $5$ and $10$ in the Weyl group $W(\mathrm{E}_{6}).$
@@
By Lemma 4.2 the order of the centralizer of an element $g$ of order $5$ in $W(\mathrm{E}_{6})$ is equal to $10$ This means that the centralizer of $g$ in the Weyl group is isomorphic either to $\mathbb{Z}/10\mathbb{Z},$ or to a dihedral group of order $10,$ since these two groups are the only groups of order $10$ up to isomorphism. However, the second case is impossible because $g$ lies in its centralizer.
@@ 5. Cubic surfaces
In this section we introduce the notation related to smooth cubic surfaces and recall the description of their geometry. More details on smooth cubic surfaces can be found in [6, Chapter 9] or [12, Chapter \@slowromancapiv@].
@@
$\Delta=\{\alpha\in\mathrm{Pic}(S_{\overline{\mathbf{F}}})\mid K_{S_{\overline{\mathbf{F}}}}\cdot\alpha=0\;\text{and}\;\alpha^{2}=-2\}.$
Then this set is the root system $\mathrm{E}_{6}.$ The Weyl group $W(\mathrm{E}_{6})$ acts on the Picard group $\mathrm{Pic}(S_{\overline{\mathbf{F}}})$ by orthogonal transformations with respect to the intersection form. This action restricts to the configuration of lines on $S_{\overline{\mathbf{F}}}.$ It fixes $K_{S_{\overline{\mathbf{F}}}}$ and preserves the cone of effective divisors.
@@
Theorem 5.1 (see, for instance, [6, Corollary 8.2.40]).
@@
Using the classification of automorphism groups of smooth cubic surfaces over an algebraically closed field we can say more about their orders.
@@
[…]
@@
This directly follows from the classification of the automorphism groups which can be found in [9, Theorem 5.3] in zero characteristic and in [7, Theorem 1.1] for positive characteristic.
@@
Lemma 5.3 ([4, Table 7.1, column 8]).
Let $S$ be a smooth cubic surface over an algebraically closed field. Consider an element $g\in W(\mathrm{E}_{6})$ of order $2$ acting on the configuration of lines on $S.$ The following assertions hold.
@@
    Assume that $g$ lies in the conjugacy class $A_{1}.$ Then there are $15$ lines on $S$ which are invariant under the action of $g.$ Moreover, under the action of $g$ there are $6$ invariant pairs of skew lines on $S,$ and in any pair the lines are interchanged by $g.$
@@
    Assume that $g$ lies in the conjugacy class $A_{1}^{2}.$ Then there are $7$ lines on $S$ which are invariant under the action of $g.$ Under the action of $g$ there are $8$ invariant pairs of skew lines on $S,$ and in any pair the lines are interchanged by $g.$ Moreover, under the action of $g$ there are $2$ invariant pairs of intersecting lines on $S,$ and in any pair the lines are interchanged by $g.$
@@
    Assume that $g$ lies in the conjugacy class $A_{1}^{3}.$ Then there are $3$ lines which are invariant under the action of $g.$ Under the action of $g$ there are $6$ invariant pairs of skew lines on $S,$ and in any pair the lines are interchanged by $g.$ Moreover, under the action of $g$ there are $6$ invariant pairs of intersecting lines on $S,$ and in any pair the lines are interchanged by $g.$
@@
    Assume that $g$ lies in the conjugacy class $A_{1}^{4}.$ Then there are $3$ lines which are invariant under the action of $g.$ Moreover, under the action of $g$ there are $12$ invariant pairs of intersecting lines on $S,$ and in any pair the lines are interchanged by $g.$
@@
By [10, Chapter 4, Proposition 2.7 and Exercise 2.9] the blowup of points in general position on a quadric surface $\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}$ gives a smooth cubic surface. In other words, the preimages of points in general position on a quadric surface are the lines in quadratic position. Conversely, the image of the blowdown of lines in quadratic position on a smooth cubic surface is a set of points in general position on $\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}.$
Let $S$ be a smooth cubic surface over a field $\mathbf{F}.$ By [12, Theorem 23.8(ii)] the Galois group $\mathrm{Gal}(\overline{\mathbf{F}}/\mathbf{F})$ acts on $\mathrm{Pic}(S_{\overline{\mathbf{F}}})$ and preserves the canonical class of $S$ and the intersection form. This means that there is a homomorphism
$\mathrm{Gal}(\overline{\mathbf{F}}/\mathbf{F})\to\mathrm{Aut}(\mathrm{Pic}(S_{\overline{\mathbf{F}}}))\subset W(\mathrm{E}_{6}).$
Let $\Gamma$ be the image of the Galois group $\mathrm{Gal}(\overline{\mathbf{F}}/\mathbf{F})$ in $W(\mathrm{E}_{6}).$ We remind the reader that $\Gamma$ lies in the centralizer of $\mathrm{Aut}(S),$ since any $\phi\in\mathrm{Aut}(S)$ is defined over $\mathbf{F}.$
Now we give auxiliary lemmas about the structures of conic bundle on del Pezzo surfaces.
@@
Let $S$ be a smooth cubic surface over an arbitrary field $\mathbf{F}.$ Assume that there is a line $L$ on $S_{\overline{\mathbf{F}}}$ which is defined over $\mathbf{F}.$ Then there is a structure of conic bundle over $\mathbb{P}^{1}$
$b\colon S\to\mathbb{P}^{1}$
on $S$ which is given by the pencil of planes on $\mathbb{P}^{3}_{\overline{\mathbf{F}}}$ passing through $L.$ This conic bundle has $5$ degenerate fibres over $\overline{\mathbf{F}}.$ Any degenerate fibre over $\overline{\mathbf{F}}$ is a union of two intersecting lines on $S_{\overline{\mathbf{F}}}.$
@@
This fact seems well-known, but for the sake of completeness let us give a proof of this fact. Let us consider the pencil of planes on $\mathbb{P}^{3}_{\overline{\mathbf{F}}}$ passing through $L.$ This pencil gives us a structure of conic bundle on $S$ over $\mathbf{F}.$ The base of this conic bundle is $\mathbb{P}^{1},$ because $L$ is defined over $\mathbf{F},$ so there is a point $p\in L$ which is defined over $\mathbf{F}.$ So the point $b(p)$ is also defined over $\mathbf{F}.$
By the very definition, any degenerate fibre over $\overline{\mathbf{F}}$ is a union of two intersecting lines on $S_{\overline{\mathbf{F}}}.$ So by computing the Euler characteristic of conic bundle we can see that there are exactly $5$ geometric degenerate fibres.
Lemma 5.7.
Let $\mathcal{F}$ be a blowup of $\mathbb{P}^{2}$ in a point. Then there is a unique structure of conic bundle on $\mathcal{F}.$
Proof.
This immediately follows from the structure of Mori cone of $\mathcal{F}.$
Now we begin studying actions of order $5$ and $10$ on smooth cubic surfaces.
Lemma 5.8.
Let $S$ be a smooth cubic surface over a field $\mathbf{F}$ such that all $27$ exceptional lines are defined over $\mathbf{F}.$ Let $\delta\in W(\mathrm{E}_{6})$ be an element of order $10$ acting on the configuration of lines on $S.$ Then
  1. (i)
    there are no lines on $S$ which are invariant under the action of $\delta;$
  2. (ii)
    there is a pair of lines on $S$ which is invariant under the action of $\delta,$ and, moreover, this pair consists of skew lines.
Proof.
Let $D$ be a cyclic subgroup in $W(\mathrm{E}_{6})$ generated by $\delta.$ Since the order of $D$ is $10,$ we get that the length of an orbit under the action of $D$ on the configuration of lines on $S$ can be $1,$ $2,$ $5$ or $10$ Let us prove that there are no orbits of length $1$ Assume the converse. Let $l$ be a $D$-invariant line on $S.$ Then we can blow down $l$ and get the del Pezzo surface $S^{\prime}$ of degree $4$ Let
$\pi\colon S\to S^{\prime}$
be this blowdown. By [6, Corollary 8.2.40] and [7, §3.1] we have
$\mathrm{Aut}(S^{\prime})\subseteq W(\mathrm{D}_{5})\simeq(\mathbb{Z}/2\mathbb{Z})^{4}\rtimes\mathfrak{S}_{5},$
where $W(\mathrm{D}_{5})$ is the Weyl group of the root system $\mathrm{D}_{5}.$ It is not hard to see that in $W(\mathrm{D}_{5})$ there are no elements of order $10$ Meanwhile, as $l$ is a $D$-invariant line, we get that $D$ acts on the configuration of lines on $S^{\prime},$ since $D$ preserves the intersection form on $\mathrm{Pic}(S_{\overline{\mathbf{F}}}).$ This means that $D\subset W(\mathrm{D}_{5}),$ which is a contradiction. So assertion (i) is proven.
Now let us prove the existence of an orbit of length $2$ under the action of $D$ which is represented by $2$ skew lines. The existence of an orbit of length $2$ follows from (i). Assume that this orbit consists of intersecting lines. Then there is a unique line $l$ which intersects both of these lines. But this means that $l$ is $D$-invariant which contradicts (i). So we get that the orbit of length $2$ consists of skew lines. So assertion (ii) is proven.
Lemma 5.9.
Let $S$ be a smooth cubic surface over a field $\mathbf{F}$ such that all $27$ exceptional lines are defined over $\mathbf{F}.$ Let $l_{1}$ and $l_{2}$ be two skew lines on $S.$ Then there are exactly $5$ lines on $S$ which intersect both $l_{1}$ and $l_{2}.$
Proof.
Let us consider the conic bundle
$b\colon S\to\mathbb{P}^{1},$
which is given by the pencil of planes on $\mathbb{P}^{3}_{\overline{\mathbf{F}}}$ passing through $l_{1}.$ By Lemma 5.6 there are $5$ degenerate fibres on $b.$ Since $l_{1}$ does not intersect $l_{2},$ we get that $l_{2}$ intersects any fibre of $b$ in a point. This means that for any degenerate fibre the line $l_{2}$ intersects one of the line in this fibre. Denote these lines in the degenerate fibres by $\mathcal{L}_{i}$ for $i=1,\ldots,5.$ Since $l_{1}$ intersects any of $\mathcal{L}_{i}$ for $i=1,\ldots,5$ by construction of $b,$ we get that $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{4},\mathcal{L}_{5}$ is the required quintuple of lines. The uniqueness follows from the fact that all lines which intersect $l_{1}$ lie in the degenerate fibres.
Lemma 5.10.
Let $S$ be a smooth cubic surface over a field $\mathbf{F}$ such that all $27$ exceptional lines are defined over $\mathbf{F}.$ Let $G$ be a group of order $5$ which acts on the configuration of lines on $S.$ Then there are two skew lines on $S$ which are $G$-invariant.
Proof.
The lengths of an orbit under the action of $G$ can be $1$ or $5$ Since $27$ is not a multiple of $5,$ we get that there are at least two $G$-invariant lines. Denote them by $l_{1}$ and $l_{2}.$ Assume that they intersect each other. Then there is a unique line which intersects both $l_{1}$ and $l_{2}.$ This means that this line is also $G$-invariant. Therefore, we obtain three $G$-invariant lines. Since $27-3=24$ is not a multiple of $5,$ we get that there are at least seven $G$-invariant lines. It is obvious, that there are two skew lines among these seven.
@@
By Lemma 5.10 there are two $G$-invariant skew lines on $S.$ Denote them by $l_{1}$ and $l_{2}.$ By Lemma 5.9 there are exactly $5$ lines on $S$ which intersect both $l_{1}$ and $l_{2}.$ Let us blow down this quintuple. Then we get a del Pezzo surface $S^{\prime}$ of degree $8$ Assume that it is a blowup of $\mathbb{P}^{2}$ in a point. Then consider two conic bundles on $S,$ say $b_{1}$ and $b_{2},$ such that
$b_{i}\colon S\to\mathbb{P}^{1}$
is given by the pencil of planes on $\mathbb{P}^{3}_{\overline{\mathbf{F}}}$ passing through $l_{i}$ for $i=1,2.$ The conic bundle $b_{i}$ gives the structure of $\mathbb{P}^{1}$-bundle on $S^{\prime}$ for $i=1,2.$ So we get two different $\mathbb{P}^{1}$-bundle on $S^{\prime}.$ By Lemma 5.7 this is impossible. Thus, we get that $S^{\prime}\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}.$